# Complete complex Finsler metrics and uniform equivalence of the Kobayashi metric

### v3

@@ 2 Terminology and preliminaries
[…]
Let us now recall the various notions of invariant metrics on a complex manifold. The Kobayashi pseudometric, denoted by $\mathfrak{K}$, is the infinitesimal form of the Kobayashi pseudodistance. Let us first recall the definition.
@@
$\begin{split}\mathfrak{K}_{M}(z,v)&=\inf\{|V|_{\mathcal{P}};V\in T^{1,0}\mathbb{D},\exists~{}\phi\in\text{Hol}(\mathbb{D},M)~{}\text{with}~{}\phi_{*}(V)=v\}\\ &=\inf\{|V|_{\mathbb{C}};V\in T^{1,0}_{0}\mathbb{D},\exists~{}\phi\in\text{Hol}(\mathbb{D},M)~{}\text{such that}~{}\phi(0)=x,~{}\phi_{*}(V)=v\}.\end{split}$ (2.1)
@@
Next, let us discuss the widely used Kobayashi distance. Denote by $(\mathbb{D},\mathcal{P})$ a unit disc in $\mathbb{C}$ equipped with the Poincaré metric
@@
[32, p. 233, Definition 9.12] A complex manifold $M$ is said to be Kobayashi hyperbolic if $d_{k}$ is a distance function, i.e., if $d_{k}(p,q)>0$ whenever $p\neq q\in M$.
3 Properties of Bergman metric on a bounded doamin
Suppose that $\mathcal{D}$ is a bounded domain. For any $z\in\mathcal{D}$, let $B(z,r(z))\subset\mathcal{D}\subset B(z,R(z))$ denote the interior and exterior with the center $z$ and the radius $r(z)$ and $R(z)$, respectively. Denote by $Sec_{\mathcal{D}}(z,v)$ and $Ric_{\mathcal{D}}(z,v)$ the holomorphic sectional curvature and the Ricci curvature with respect to the Bergman metric at the point $z$ in the direction $v$. Denote by $Scal_{\mathcal{D}}(z)$ the scalar curvature of the Bergman metric at the point $z$. In [17], Lu obtained some lower and upper bounds of the holomorphic sectional curvature and Ricci curvature of the Bergman metric on $\mathcal{D}$. Not long after, Zhang [31] obtained the estimates of the scalar curvature of the Bergman metric of $\mathcal{D}$. The methods rely on constructing proper minimal functions. However, we note that some lower bounds may tends to $\infty$- ∞ at the point $z$ goes to the boundary $\partial D$. In what follows, we can see that there are many bounded domains, with the so-called uniformly squeezing property, whose Bergman curvature admits a uniform lower bound [31].
@@ 3 Properties of Bergman metric on a bounded doamin
[19, p. 629, Definition 7.2][30, p. 548, Definition 1] A complex manifold $M$ of complex dimension $n$ is called uniformly squeezing (USq) or equivalently holomorphic homogeneous regular (HHR), if there are uniform positive constants $r$ and $R$ such that for any point $p\in M$, there is a holomorphic map $f_{p}:M\rightarrow\mathbb{C}^{n}$ which satisfies
@@
iii). $B^{n}(0,r)\subset f_{p}(M)\subset B^{n}(0,R)$, where $B^{n}(0,r)$ and $B^{n}(0,R)$ are Euclidean balls with center 00 in $\mathbb{C}^{n}$.
@@
where the supremum is taken over all holomorphic embeddings $f:\mathcal{D}\rightarrow B^{n}$ with $f(z)=0$, $B^{n}$ is the unit ball in $\mathbb{C}^{n}$, and $B^{n}(0,r)$ is the ball in $\mathbb{C}^{n}$ with center 00 and radius $r$. As $z$ varies, we get a function $s_{\mathcal{D}}$ on $\mathcal{D}$, which is called the squeezing function of $\mathcal{D}$.
@@
[31, p. 1150, Theorem 1.1] For any point $z\in\mathcal{D}$ and $v\in T^{1,0}_{z}\mathcal{D}\backslash\{0\}\cong\mathbb{C}^{n}\backslash\{0\}$, then
$2-2\frac{n+2}{n+1}s_{\mathcal{D}}^{-4n}\leq Sec_{\mathcal{D}}(z,v)\leq 2-2\frac{n+2}{n+1}s_{\mathcal{D}}^{4n},$ (3.1)
$(n+1)-2(n+2)s_{\mathcal{D}}^{-2n}\leq Ric_{\mathcal{D}}(z,v)\leq(n+1)-(n+2)s_{\mathcal{D}}^{2n},$ (3.2)
$n(n+1)-n(n+2)s_{\mathcal{D}}^{-2n}\leq Scal_{\mathcal{D}}(z)\leq n(n+1)-n(n+2)s_{\mathcal{D}}^{2n}.$ (3.3)
In [8], Deng et al. proved that any bounded strongly pseudoconvex domains with $C^{2}$-boundary in $\mathbb{C}^{n}$ are HHR. In [11], Kim and Zhang showed that all bounded convex domains in $\mathbb{C}^{n}$ are HHR. Combining Theorem 3.1, the above results and Remark 3.1, we obtain the following theorem.
Theorem 3.2.
Suppose that $\mathcal{D}$ is a bounded strictly pseudoconvex domain with $C^{2}$-boundary or bounded convex domain in $\mathbb{C}^{n}$. For each point $z\in\mathcal{D}$ and $v\in T^{1,0}_{z}\mathcal{D}\backslash\{0\}\cong\mathbb{C}^{n}\backslash\{0\}$, then
$C_{1}\leq Sec_{\mathcal{D}}(z,v)\leq C_{2},$ (3.4)
$C_{3}\leq Ric_{\mathcal{D}}(z,v)\leq C_{4},$ (3.5)
$C_{5}\leq Scal_{\mathcal{D}}(z)\leq C_{6},$ (3.6)
where $C_{1},C_{2},C_{3},C_{4},C_{5},C_{6}$ are only dependent in complex dimension $n$ of the complex manifold $\mathcal{D}$, $C_{1},C_{3},C_{5}$ are negative constants, $C_{2},C_{4},C_{6}$ are positive constants.
@@
[30, p. 550, Corollary 3] Let $\mathcal{D}$ be a bounded domain with the uniform squeezing property. Denote by $d=d(z,\partial\mathcal{D})$ the Euclidean distance of $z\in\mathcal{D}$ from the boundary $\partial\mathcal{D}$ of $\mathcal{D}$. Then $K(z,z)\geq\frac{c}{d^{2}(-\log d)^{2}}$ for some constant $c$.
@@
4 Holomorphic sectional curvature and sectional curvature on a Kähler manifold
In this section, we introduce some relationship between holomorphic sectional curvature and sectional curvature. In order to apply Schwarz lemma from a complete Kähler manifold into a complex Finsler manifold, we need to know relationship between the bound of holomorphic sectional curvature and sectional curvature in a Kähler manifold. In Kähler geometry, we can easily know that holomorphic sectional curvature is ’dominated’ by the sectional curvature [32]. Therefore, bound of holomorphic sectional curvature can be derived from bound of sectional curvature. It is very natural to ask the question: If holomorphic sectional curvature of a Kähler manifold is bounded , can we deduce that its sectional curvature is bounded? Now, we give affirmative answer to above question. Firstly, we can refer to an exercise in [32, p. 189, Exercise 17]. In this paper, we don’t need such a strong theorem. Hence, we introduce and prove the following theorem.
@@ 4 Holomorphic sectional curvature and sectional curvature on a Kähler manifold
From properties of curvature of Kähler manifolds, we obtain
@@
5 Some intrinsic metrics
For a complex manifold, there are natural and intrinsic complex Finsler pseudometric, i.e., the Kobayashi pseudometric and the Carathéodory pseudometric. It is known that the Kobayashi pseudometric is the maximum pseudometric among pseudometric satisfying the decreasing property. However, the Carathéodory pseudometric is the minimum. If we consider bounded domains in $\mathbb{C}^{n}$, both Kobayashi pseudometric and Carathéodory pseudometric are non-degenerate complex Finsler metrics. In 1980s, Lempert [15, 16] proved the following theorem.
Theorem 5.1.
$\mathfrak{K}_{\mathcal{D}}=\mathfrak{C}_{\mathcal{D}}.$
Remark 5.1.
[16, see p. 341-343] As long as $\mathcal{D}$ is a bounded strongly convex domain with smooth boundary in $\mathbb{C}^{n}$, the differentiability condition of two intrinsic metrics are satisfied in sense of Abate and Patrizio in this paper.
In [27], Wong proved the holomorphic sectional curvature of the Kobayashi metric $\mathfrak{K}_{M}$ is greater than or equal to $-4$. Aslo, he proved the holomorphic sectional curvature of the Carathéodory metric $\mathfrak{C}_{M}$ is less than or equal to $-4$. By Theorem 5.1, we know that the Kobayashi metric on a bounded strongly convex domain with smooth boundary is complex Finsler metric in sense of Abate and Patrizio whose holomorphic sectional curvature is constant and equal to $-4$.
Definition 5.1.
$a)$ it is smooth function on $\widetilde{M}$ with values $\mathbb{R}^{+}$;
$(3)$ the metric, which is induced by $\mathfrak{K}_{M}$ on the totally geodesic complex curve $\gamma_{\omega}(U)$, is Kähler and with constant holomorphic sectional curvature equal to $-4$;
Remark 5.2.
[25, p. 292, Theorem 1.3] In 1990, Faran proved a complex manifold $M$ is Lempert manifold if and only if it admits a strongly pseudoconvex complex Finsler metric $G$, which verifies $(2),(3)$, and $(4)$ of Definition 5.1. In this case $F$ concides with the Kobayashi metric $\mathfrak{K}_{M}$. By Proposition 2.1, we know that the holomorphic sectional curvature of Lempert manifolds is equal to $-4$.
An immediate interest for Lempert manifolds come from the well-known result of Lempert on the Kobayashi metric of strongly convex domians in $\mathbb{C}^{n}$.
Theorem 5.2.
@@ 6 Strongly negative holomorphic sectional curvature of complex Finsler and Kobayashi-hyperbolic
In [33], Zhong find that there are many explicit non-Hermitian quadratic strongly pseudoconvex and convex complex Finsler metrics on a domain $\mathcal{D}$ in $\mathbb{C}^{n}$. Inquisitively, what are the holomorphic sectional curvature of these complex Finsler metrics? It is a interesting question. In [21], Nie and Zhong proved a bounded domain admits a non-Hermitian quadratic strongly pseudoconvex complex Finsler metric $G:T^{1,0}\mathcal{D}\rightarrow[0,+\infty)$ such that its holomorphic sectional curvature is bounded from above by a negative constant. It is very natural to ask whether these complex Finsler metrics are complete? It is known that any bounded domains in $\mathbb{C}^{n}$ are Kobayashi hyperbolic. We wonder what bounded domain is complete Kobayashi hyperbolic. In this section, we answer these questions. Before giving the proofs of our main results, we need some results as follows. The first one is the Schwarz lemma from a complete Kähler manifold into a complex Finsler manifold.
@@
[…]
$G(z,v):=r\phi(t,s)\quad\mbox{with}\quad\phi(t,s)=e^{at+bs},\quad\forall z\in\mathcal{D},v\in T_{z}^{1,0}\mathcal{D}.$ (6.2)
It is clear that $G:T^{1,0}\mathcal{D}\rightarrow[0,+\infty)$ is a non-Hermitian quadratic strongly pseudoconvex complex Finsler metric such that its holomorphic sectional curvature is bounded from above by a negative constant from Theorem 6.4 in [21] and Proposition 2.9 in [33]. If $b$ is equal to 00 and $a$ is greater than 00, $G$ is a Hermitian metric such that its holomorphic sectional curvature is bounded from above by a negative constant.
Theorem 6.3.
[13, p. 233, Theorem 4.10.24] [30, p. 548, Theorem 1] Every bounded pseudoconvex domain in $\mathbb{C}^{n}$ with a $C^{1}$-boundary or domain with the uniform squeezing property is complete with respect to its Bergman metric.
Theorem 6.4.
For a suitable choice of positive constant $C$, suppose that $\mathcal{D}$ is a bounded strictly pseudoconvex domain with $C^{2}$-boundary or bounded convex domain in $\mathbb{C}^{n}$. Suppose that $\mathcal{D}$ admits a strongly pseudoconvex complex Finsler metric $G:T^{1,0}\mathcal{D}\rightarrow[0,+\infty)$ such that its holomorphic sectional curvature is bounded from above by a negative constant $-K$. Then $H=CG+g_{B}$ and the Bergman metric $g_{B}$ are equivalent. What’s more, $H=CG+g_{B}$ is complete.
Proof.
$\forall v\in T^{1,0}_{z}\mathcal{D}\backslash\{0\}\cong\mathbb{C}^{n}\backslash\{0\}$. By assumptions and Theorem 3.2, we observed that the holomorphic sectional curvature $Sec_{\mathcal{D}}(z,v)$ along a non-zero tangent vector $v$ of the Bergman metric $g_{B}$ on $\mathcal{D}$ is bounded. We have
$-A\leq Sec_{\mathcal{D}}(z,v)\leq B,$ (6.3)
where $A,B$ are positive constants.
By the equality $H=CG+g_{B}$ and Proposition 2.1, we have
$\begin{split}K_{H}(v)=\sup_{\varphi}\{K(\varphi^{*}H)(0)\}&=\sup_{\varphi}\{K(\varphi^{*}(CG)+\varphi^{*}g_{B})(0)\}\\ &=\sup_{\varphi}\{\frac{1}{C}K(\varphi^{*}G)(0)+K(\varphi^{*}g_{B})(0)\}\\ &\leq\frac{1}{C}\sup_{\varphi}\{K(\varphi^{*}G)(0)\}+\sup_{\varphi}\{K(\varphi^{*}g_{B})(0)\}\\ &=\frac{1}{C}K_{G}(v)+Sec_{\mathcal{D}}(z,v),\end{split}$ (6.4)
where $K_{H}(v),K_{G}(v)$ are the holomorphic sectional curvature along a non-zero tangent vector $v$ of the complex Finsler metrics $H,G$ respectively. Here the supremum is taken with respect to the family of all holomorphic maps $\varphi:\mathbb{D}\rightarrow M$ with $\varphi(0)=z$ and $\varphi^{\prime}(0)=\lambda v$ for some $\0\lambda\in\mathbb{C}\backslash\{0\}$.
@@
We take $0<C<\frac{K}{B}$, we know that the holomorphic sectional curvature $K_{H}(v)$ is bounded from above by a negative constant $-\frac{K}{C}+B$.
Combining with Theorem 4.1, we know that the sectional curvature of the Bergman metric $g_{B}$ on $\mathcal{D}$ is also bounded. By Theorem 3.2, we know that the holomorphic sectional curvature of the Bergman metric on $\mathcal{D}$ is bounded from below by a negative constant $-C_{1}$. Applying Theorem 6.1 to the case where $f$ is an identity holomorphic map $id$, we have
@@
[13, p. 112, Theorem 3.7.1 ] Suppose that $(M,G)$ is a complete strongly pseudoconvex complex Finsler manifold whose holomorphic sectional curvature is bounded from above by a negative constant $-B<0$. Then $M$ is a complete Kobayashi-hyperbolic.
@@
$G(z,v)=(\phi^{*}G)(0,\xi)\leq\frac{4}{B}\omega_{\mathcal{P}}(0,\xi)=\frac{4}{B}|\xi|^{2}_{\mathbb{C}}.$ (6.8)
@@
As a simple application of Theorems 6.2, 6.4, 6.5 and Remark 6.1, we obtain the following corollary.

### v4

@@ 2.2 Invariant metrics
where $\|\cdot\|$ stands for the $L^{2}$-norm. We may assume that $\sup_{f\in H_{(2)}^{0}\left(M,K_{M}\right),\|f\|=1}\left|f_{U}(x)\right|$ is realized by $f_{x}\in H_{(2)}^{0}(M,K)$ with $\left\|f_{x}\right\|=1$ so that $K_{U}(x,x)=\left|f_{x,U}(x)\right|^{2}$.
For a complex manifold, there are natural and intrinsic complex Finsler pseudometric, i.e., the Kobayashi pseudometric and the Carathéodory pseudometric. It is known that the Kobayashi pseudometric is the maximum pseudometric among pseudometric satisfying the decreasing property. However, the Carathéodory pseudometric is the minimum. Let us recall some definitions.
@@
$\begin{split}\mathfrak{K}_{M}(z,v)&=\inf\{|V|_{\mathcal{P}};V\in T^{1,0}\mathbb{D},\exists~\phi\in\text{Hol}(\mathbb{D},M)~\text{with}~\phi_{*}(V)=v\}\\ &=\inf\{|V|_{\mathbb{C}};V\in T^{1,0}_{0}\mathbb{D},\exists~\phi\in\text{Hol}(\mathbb{D},M)~\text{such that}~\phi(0)=x,~\phi_{*}(V)=v\}.\end{split}$
@@
If we consider bounded domains in $\mathbb{C}^{n}$, both Kobayashi pseudometric and Carathéodory pseudometric are non-degenerate complex Finsler metrics. In 1980s, Lempert [15, 16] proved the following theorem.
Theorem 2.1.
$\mathfrak{K}_{\mathcal{D}}=\mathfrak{C}_{\mathcal{D}}.$
Remark 2.1.
[16, see p. 341-343] As long as $\mathcal{D}$ is a bounded strongly convex domain with smooth boundary in $\mathbb{C}^{n}$, the differentiability condition of two intrinsic metrics is satisfied in sense of Abate and Patrizio in this paper.
In [28], Wong proved the holomorphic sectional curvature of the Kobayashi metric $\mathfrak{K}_{M}$ is greater than or equal to 4-4. Aslo, he proved the holomorphic sectional curvature of the Carathéodory metric $\mathfrak{C}_{M}$ is less than or equal to 4-4. By Theorem 2.1, we know that the Kobayashi metric on a bounded strongly convex domain with smooth boundary is complex Finsler metric in sense of Abate and Patrizio whose holomorphic sectional curvature is constant and equal to 4-4.
Definition 2.4.
$a)$ it is smooth function on $\widetilde{M}$ with values $\mathbb{R}^{+}$;
$(3)$ the metric, which is induced by $\mathfrak{K}_{M}$ on the totally geodesic complex curve $\gamma_{\omega}(U)$, is Kähler and with constant holomorphic sectional curvature equal to 4-4;
Remark 2.2.
[25, p. 292, Theorem 1.3] In 1990, Faran[9] proved a complex manifold $M$ is Lempert manifold if and only if it admits a strongly pseudoconvex complex Finsler metric $G$, which verifies $(2),(3)$, and $(4)$ of Definition 2.4. In this case $F$ coincides with the Kobayashi metric $\mathfrak{K}_{M}$. By Proposition 2.3, we know that the holomorphic sectional curvature of Lempert manifolds is equal to 4-4.
An immediate interest for Lempert manifolds comes from the well-known result of Lempert on the Kobayashi metric of strongly convex domains in $\mathbb{C}^{n}$.
Theorem 2.2.
2.3 Kobayashi invariant distance
Now let us discuss the widely used Kobayashi distance. Denote by $(\mathbb{D},\mathcal{P})$ a unit disc in $\mathbb{C}$ equipped with the Poincaré metric
@@ 2.3 Kobayashi invariant distance
[34, p. 233, Definition 9.12] A complex manifold $M$ is said to be Kobayashi hyperbolic if $d_{k}$ is a distance function, i.e., if $d_{k}(p,q)>0$ whenever $p\neq q\in M$.
2.4 Curvature of Bergman metric on a bounded domain
Suppose that $\mathcal{D}$ is a bounded domain. For any $z\in\mathcal{D}$, let $B(z,r(z))\subset\mathcal{D}\subset B(z,R(z))$ denote the interior and exterior with the center $z$ and the radius $r(z)$ and $R(z)$, respectively. Denote by $Sec_{\mathcal{D}}(z,v)$ and $Ric_{\mathcal{D}}(z,v)$ the holomorphic sectional curvature and the Ricci curvature with respect to the Bergman metric at the point $z$ in the direction $v$. Denote by $Scal_{\mathcal{D}}(z)$ the scalar curvature of the Bergman metric at the point $z$.
@@ 2.4 Curvature of Bergman metric on a bounded domain
[19, p. 629, Definition 7.2][32, p. 548, Definition 1] A complex manifold $M$ of complex dimension $n$ is called uniformly squeezing (USq) or equivalently holomorphic homogeneous regular (HHR), if there are uniform positive constants $r$ and $R$ such that for any point $p\in M$, there is a holomorphic map $f_{p}:M\rightarrow\mathbb{C}^{n}$ which satisfies
@@
iii). $B^{n}(0,r)\subset f_{p}(M)\subset B^{n}(0,R)$, where $B^{n}(0,r)$ and $B^{n}(0,R)$ are Euclidean balls with center 0 in $\mathbb{C}^{n}$.
@@
where the supremum is taken over all holomorphic embeddings $f:\mathcal{D}\rightarrow B^{n}$ with $f(z)=0$, $B^{n}$ is the unit ball in $\mathbb{C}^{n}$, and $B^{n}(0,r)$ is the ball in $\mathbb{C}^{n}$ with center 0 and radius $r$. As $z$ varies, we get a function $s_{\mathcal{D}}$ on $\mathcal{D}$, which is called the squeezing function of $\mathcal{D}$.
@@
[33, p. 1150, Theorem 1.1] For any point $z\in\mathcal{D}$ and $v\in T^{1,0}_{z}\mathcal{D}\backslash\{0\}\cong\mathbb{C}^{n}\backslash\{0\}$, then
$2-2\frac{n+2}{n+1}s_{\mathcal{D}}^{-4n}\leq Sec_{\mathcal{D}}(z,v)\leq 2-2\frac{n+2}{n+1}s_{\mathcal{D}}^{4n},$
$(n+1)-2(n+2)s_{\mathcal{D}}^{-2n}\leq Ric_{\mathcal{D}}(z,v)\leq(n+1)-(n+2)s_{\mathcal{D}}^{2n},$
$n(n+1)-n(n+2)s_{\mathcal{D}}^{-2n}\leq Scal_{\mathcal{D}}(z)\leq n(n+1)-n(n+2)s_{\mathcal{D}}^{2n}.$
@@
[32, p. 550, Corollary 3] Let $\mathcal{D}$ be a bounded domain with the uniform squeezing property. Denote by $d=d(z,\partial\mathcal{D})$ the Euclidean distance of $z\in\mathcal{D}$ from the boundary $\partial\mathcal{D}$ of $\mathcal{D}$. Then $K(z,z)\geq\frac{c}{d^{2}(-\log d)^{2}}$ for some constant $c$.
@@
2.5 Holomorphic sectional curvature and sectional curvature on a Kähler manifold
In this section, we introduce some relationship between holomorphic sectional curvature and sectional curvature. In Kähler geometry, we can easily know that holomorphic sectional curvature is ’dominated’ by the sectional curvature [34]. It is very natural to ask the question: If holomorphic sectional curvature of a Kähler manifold is bounded , can we deduce that its sectional curvature is bounded? Now, we give affirmative answer to above question. Firstly, we can refer to an exercise in [34, p. 189, Exercise 17]. In this paper, we don’t need such a strong theorem. Hence, we introduce and prove the following theorem.
@@ 2.5 Holomorphic sectional curvature and sectional curvature on a Kähler manifold
From properties of curvature of the Kähler manifold $(M,h)$, we obtain
@@ 3 Strongly negative holomorphic sectional curvature of complex Finsler and Kobayashi-hyperbolic
In this section, we will prove the Theorem 1.1 (i.e. Theorem 3.4) and 1.2 (i.e. Corollary 3.1). Before giving the proofs of our main results, we need some results as follows. The first one is the Schwarz lemma from a complete Kähler manifold into a complex Finsler manifold.
@@
here $\|\cdot\|$ denotes the canonical complex Euclidean norm of $z\in\mathcal{D}$. For any $z\in\mathcal{D},0\neq v\in T_{z}^{1,0}\mathcal{D}=\mathbb{C}^{n}$, we denote
where $\langle\cdot,\cdot\rangle$ denotes the canonical complex Euclidean inner product in $\mathbb{C}^{n}$. For every constants $a,b$ satisfying $a>0$ and $0<b<\frac{1}{M_{0}}$, we define
$G(z,v):=r\phi(t,s)\quad\mbox{with}\quad\phi(t,s)=e^{at+bs},\quad\forall z\in\mathcal{D},v\in T_{z}^{1,0}\mathcal{D}.$
It is clear that $G:T^{1,0}\mathcal{D}\rightarrow[0,+\infty)$ is a non-Hermitian quadratic strongly pseudoconvex complex Finsler metric such that its holomorphic sectional curvature is bounded from above by a negative constant from Theorem 6.4 in [21] and Proposition 2.9 in [35]. If $b$ is equal to 0 and $a$ is greater than 0, $G$ is a Hermitian metric such that its holomorphic sectional curvature is bounded from above by a negative constant.
In [8], Deng et al. proved that any bounded strictly pseudoconvex domains with $C^{2}$-boundary in $\mathbb{C}^{n}$ are HHR. In [11], Kim and Zhang showed that all bounded convex domains in $\mathbb{C}^{n}$ are HHR. Combining Theorem 2.3, the above results and Remark 2.3, we obtain the following Lemma.
Lemma 3.1.
Suppose that $\mathcal{D}$ is a bounded strictly pseudoconvex domain with $C^{2}$-boundary or bounded convex domain in $\mathbb{C}^{n}$. Then holomorphic sectional curvature, Ricci curvature, scalar curvature of the Bergman metric on $\mathcal{D}$ are bounded.
Theorem 3.3.
[24, p. 238, Theorem 1] [32, p. 548, Theorem 1] Every bounded pseudoconvex domain in $\mathbb{C}^{n}$ with a $C^{1}$-boundary or domain with the uniform squeezing property is complete with respect to its Bergman metric.
Theorem 3.4.
For a suitable choice of positive constant $C$, suppose that $\mathcal{D}$ is a bounded strictly pseudoconvex domain with $C^{2}$-boundary or bounded convex domain in $\mathbb{C}^{n}$. Suppose that $\mathcal{D}$ admits a strongly pseudoconvex complex Finsler metric $G:T^{1,0}\mathcal{D}\rightarrow[0,+\infty)$ such that its holomorphic sectional curvature is bounded from above by a negative constant K-K. Then $H=CG+g_{B}$ and the Bergman metric $g_{B}$ are equivalent. What’s more, $H=CG+g_{B}$ is complete.
Proof.
$\forall v\in T^{1,0}_{z}\mathcal{D}\backslash\{0\}\cong\mathbb{C}^{n}\backslash\{0\}$. By assumptions, we observed that the holomorphic sectional curvature $Sec_{\mathcal{D}}(z,v)$ along a non-zero tangent vector $v$ of the Bergman metric $g_{B}$ on $\mathcal{D}$ is bounded. By Lemma 3.1, we have
$A\leq Sec_{\mathcal{D}}(z,v)\leq B,$ (3.1)
where $A,B$ are positive constants.
Combine the equality $H=CG+g_{B}$ with Propositions 2.1, 2.2, 2.3 , we have
$\begin{split}K_{H}(v)=\sup_{\varphi}\{K(\varphi^{*}H)(0)\}&=\sup_{\varphi}\{K(\varphi^{*}(CG)+\varphi^{*}g_{B})(0)\}\\ &\leq\sup_{\varphi}\{\frac{1}{C}K(\varphi^{*}G)(0)+K(\varphi^{*}g_{B})(0)\}\\ &\leq\frac{1}{C}\sup_{\varphi}\{K(\varphi^{*}G)(0)\}+\sup_{\varphi}\{K(\varphi^{*}g_{B})(0)\}\\ &=\frac{1}{C}K_{G}(v)+Sec_{\mathcal{D}}(z,v),\end{split}$ (3.2)
@@
We take $0<C<\frac{K}{B}$, we know that the holomorphic sectional curvature $K_{H}(v)$ is bounded from above by a negative constant $\frac{K}{C}+B$.
Combining with Theorem 2.5, we know that the sectional curvature of the Bergman metric $g_{B}$ on $\mathcal{D}$ is also bounded. By lemma 3.1, we know that the holomorphic sectional curvature of the Bergman metric on $\mathcal{D}$ is bounded from below by a negative constant $C_{1}$. Applying Theorem 3.1 to the case where $f$ is an identity holomorphic map $id$, we have
@@
[13, p. 112, Theorem 3.7.1 ] Suppose that $(M,G)$ is a complete strongly pseudoconvex complex Finsler manifold whose holomorphic sectional curvature is bounded from above by a negative constant B<0-B<0. Then $M$ is a complete Kobayashi-hyperbolic.
@@
$G(z,v)=(\phi^{*}G)(0,\xi)\leq\frac{4}{B}\omega_{\mathcal{P}}(0,\xi)=\frac{4}{B}|\xi|^{2}_{\mathbb{C}}.$
@@
As a simple application of Theorems 3.2, 3.4, 3.5 and Remark 3.1, we obtain the following corollary.